OpenStudy (anonymous):

Cam anyone help me? 2- 1 1/2 dozen=? Cam anyone help me? 2- 1 1/2 dozen=? @Mathematics

hero (hero):

Can or Cam?

OpenStudy (karatechopper):

whats the first 2-11?

OpenStudy (anonymous):

Can sorry

OpenStudy (karatechopper):

hero its a mistake let it go

hero (hero):

I like pointing out mistakes. It's a bad habit of mine. Sorry

OpenStudy (anonymous):

and the question is Take away 2 from 1 1/2 dozen! ahaha

hero (hero):

:P

hero (hero):

2 - 1.5 = ?

OpenStudy (anonymous):

See this is y problem, I type and it looks right then watcha know its not right! and yes hero

hero (hero):

1/2 dozen which is actually 6

OpenStudy (anonymous):

omg MY*

OpenStudy (anonymous):

1 1/2

OpenStudy (anonymous):

so i get it!! of so a dozen is 12 cause 6+6=12 half of 12 is 6!!

hero (hero):

You have to do the math first, then translate it afterwards.

OpenStudy (anonymous):

haaha! But 1 dozen is 12

OpenStudy (anonymous):

so what is half?

hero (hero):

1/2 dozen = 6 dozen. I thought you already had this figured out

OpenStudy (anonymous):

cause 1 and a 1/2... and how do you take away 2? when there is only 1 1/2

OpenStudy (anonymous):

Oh yeah... but ^_^

OpenStudy (karatechopper):

so 2 dozen is 24

OpenStudy (anonymous):

so 24 or 12 >.< this is why I hate math XD

OpenStudy (anonymous):

cuz 2 dozen is 24 and 1 is 12 so? half of 12 is 6 and half of 24 is 12..

hero (hero):

24 - 18 = 6

hero (hero):

2 dozen = 24 1.5 dozen = 18

OpenStudy (anonymous):

So.. 6? :D

hero (hero):

Yes

OpenStudy (anonymous):

ok thanks